Problem 9-3 Lower of cost or market [LO9-1 Forester Company has five products in its inventory. Information about the December 31, 2018, inventory follows. Unit Unit Unit Replaceent Selling Product Quantity Cost s 17 Cost $19 Price $23 25 1,000 700 1,000 900 600 18 10 14 21 15 13 19 20 The cost to sell for each product consists of a 15 percent sales commission. The normal profit percentage for each product is 40 percent of the selling price Required: 1. Determine the carrying value of inventory at December 31, 2018, assuming the lower of cost or market (LCM) rule is applied to individual products. 2a. Determine the carrying value of inventory at December 31, 2018, assuming the LCM rule is applied to the entire inventory. 2b. Assuming inventory write-downs are usual business practice for Forester, record any necessary year-end adjusting entry.
Complete this question by entering your answers in the tabs below. Req 1 Req 2AReq 28 Determine the carrying value of inventory at December 31, 2018, assuming the lower of cost or market (LCM) rule is applied to individual products. (Do not round intermediate calculations.) Product units) A (1000) B (700) C (1000) D (900) E (600) nventory Value RC NRV NRV-NP Market Cost Total
Complete this question by entering your answers in the tabs below. Req 1 Req 2AReq 2B Determine the carrying value of inventory at December 31, 2018, assuming the LCM rule is applied to the entire inventory. (Do not round intermediate calculations.) Inventory carrying value K Req 1
Req 1 Req 2A Rq 28 Assuming the inventory write-downs are usual business practice for Forester, record any necessary year-end adjusting entry. (Do not round intermediate calculations. If no entry is required for a transaction/event, select "No journal entry required" in the first account field.) View transaction list Journal entry worksheet Record any necessary year-end adjusting entry assuming that inventory write-downs are common for Forester Company Note: Enter debits before credits. Event General Journal Debit Credit
Answer #1
Inventory value with LCM applied to entire Inventory
A B C D E=A*B F=A*C G=A*D
Unit Unit Total Total
Product Quantity Unit Cost Replacement Cost Selling Price Total Cost Replacement Cost Market value
A          1,000 $17 $19 $23 $17,000 $19,000 $23,000
B             700 $22 $18 $25 $15,400 $12,600 $17,500
C          1,000 $10 $9 $15 $10,000 $9,000 $15,000
D             900 $14 $11 $13 $12,600 $9,900 $11,700
E             600 $21 $19 $20 $12,600 $11,400 $12,000
TOTAL $67,600 $61,900 $79,200
Cost to sell $11,880 (79200*0.15)
(market Ceiling) Net Realizable Value(NRV) $67,320 (79200-11880)
Normal Profit $31,680 (0.4*79200)
Market floor $35,640
Replacement Cost $61,900
Market $61,900
NRV=Market Value -Cost To Sell Historical Cost $67,600
NRV=Market Value *(1-0.15) Lower of Cost or market $61,900 (LCM)
NP=Net Profit =0.4*Market value
F H=G*(1-0.15) I=H-0.4*G E
Product Quantity RC NRV NRV-NP Market Cost Inventory Value
A          1,000 $19,000 $19,550 $10,350 $19,000 $17,000 $17,000
B             700 $12,600 $14,875 $7,875 $12,600 $15,400 $12,600
C          1,000 $9,000 $12,750 $6,750 $9,000 $10,000 $9,000
D             900 $9,900 $9,945 $5,265 $9,900 $12,600 $9,900
E             600 $11,400 $10,200 $5,400 $10,200 $12,600 $10,200
Total $61,900 $67,320 $35,640 Total Cost $67,600 $58,700 (LCM)
Inventory value with LCM applied to individual Product $58,700
Inventory value with LCM applied to entire Inventory $61,900
JOURNAL ENTRY
Account Debit Credit
Loss on LCM adjustment $8,900 (67600-58700)
Finished goods Inventory $8,900
LCM applied to entire inventory:
Loss on LCM adjustment $5,700 (67600-61900)
Finished goods Inventory $5,700